OpenStudy (anonymous):

i have two points (1,5) and (3,1) im trying to find my 3rd point but dont know how....and how do i get my slope with the given info

OpenStudy (anonymous):

the slope = (1-5) / (3-1) = -4/2 = -2

OpenStudy (anonymous):

so i just subtract the points??? to get the slope y=-2x+

OpenStudy (anonymous):

how do i get the y and x intercept

OpenStudy (anonymous):

the equation of the line will be y-1 = -2(x-3) y = -2x + 7 to find a third point just plug in a value for x eg let x = 2 then y = -2(2) + 7 = 3 this point is (2,3)

OpenStudy (anonymous):

y intercept is value of y when x = 0 so its y = 0 +7 = 7

OpenStudy (anonymous):

x intercept is value of x when y = 0 so 0 = -2x + 7 2x = 7 x = 3.5 = x intercept

OpenStudy (anonymous):

so to find a third point what equation do i use

OpenStudy (anonymous):

u use y= -2x + 7 pick a value for x and plug it into this equation eg let x = 4 then y = -2*4 + 7 = -8 + 7 = -1 so a third point is (4,-1)

OpenStudy (anonymous):

the line look something like:

OpenStudy (anonymous):